OpenStudy (anonymous):

radical 54 minus radical 1/6? what now? can somebodyhelpme??? lol

OpenStudy (saifoo.khan):

\[\Huge \text{ try it out!!!}\]

OpenStudy (anonymous):

ive been trying while i was waiting on ppl to help me with other probs ive been working these out and FAILING!! D: MY BRAIN NO LIKE FRACTIONS!! LOL im dead frickin serious here.

OpenStudy (saifoo.khan):

ATleast try.

OpenStudy (paxpolaris):

\[\sqrt{54}-\sqrt{\frac16}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I DID TRY.

OpenStudy (paxpolaris):

first rationalize the denominator of the fraction...

OpenStudy (anonymous):

whocanawhatchit? im a decimal person, me no like fractions. me no compute. help. lol

OpenStudy (anonymous):

just a small hint : \[\sqrt{1/6} = \sqrt{1}/\sqrt{6}\]

OpenStudy (saifoo.khan):

\[\sqrt{54} + \frac{\sqrt1}{\sqrt6}\]

OpenStudy (anonymous):

i am banging my head against the wall repeatedly. i am not good with fractions!

OpenStudy (anonymous):

Take LCM

OpenStudy (saifoo.khan):

\[\Facepalm\]

OpenStudy (anonymous):

\[3\sqrt{6}+1/6\sqrt{6}=(19\sqrt{6})6\]

OpenStudy (anonymous):

over 6*

OpenStudy (paxpolaris):

\[\sqrt \frac16 =\frac {\sqrt1} {\sqrt6} = \frac 1 {\sqrt6}\]

OpenStudy (anonymous):

ok i think i got it thanks yall! :)

OpenStudy (saifoo.khan):

Finally! ;)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[{\sqrt{54}}-{{1}\over{\sqrt{6}}}\]

OpenStudy (anonymous):

now do the simple calculation

OpenStudy (paxpolaris):

\[\huge {17 \sqrt6 \over 6}\]

OpenStudy (anonymous):

thanks i got it right lol :)

OpenStudy (anonymous):

sorry i was adding